OpenStudy (anonymous):

Solve k = 1/3 (b-2a) for b. Show work please!

OpenStudy (saifoo.khan):

\[b = 2 a+3 k \]

OpenStudy (anonymous):

so whats the solution to that^

OpenStudy (anonymous):

k= 1/3 (b - 2a) 3k = b -2a 2a + 3k = b

OpenStudy (saifoo.khan):

yes that's the solution.